Controls Meeting 9-Aug-2012
- FPGA programming tutorial by Ben Raydo some time in Aug at Online meeting
- Beni will talk about the Gas System in early Sep.
- Followup on Danfysik supplies by Sarin 23-Aug
- Minutes from previous meeting 26-Jul-2012
- Hardware purchase status - Josh, Hovanes
- EPICS - Hovanes, Vanik
- motion control, alarm system, goniometer, other power supplies, etc.
- When will Nerses arrive?
- Counting house installation - Hovanes
- Solenoid test - Mark, Josh, Scot, Elliott
- Fasttrack schedule
- Power supply maintenance OSP
- Goals, test plan, procedures, FMEA, reviews, operations OSP
- Vacuum, SOE, quench detector, etc.
- PLC programming and PXI system - Josh, Dave, Yi
- Test stand - Hovanes, Josh
- JInventory improvements - Elliott
- New Accelerator Elog - Elliott
- Mantis task tracker
- Midas elog
- Experiment Control Supervisor - Elliott
NOTE change of time
10AM Thurs 26-Jul-2012 CC L207
9:30am Thurs 23-Aug-2012 CC L207
Present: Elliott W, Mark S, Hovanes E, Josh B, Vanik K.
The stuff Mark, Hovanes and Josh ordered is coming in: tools, electronics modules, computers and monitors, etc.
Vanik leaves in about a week and is scheduled to return in Jan to work on the goniometer. He completed modifications to the XPS EPICS driver, supporting all features we need, including pulse output to drive DAQ in a Harp scan. Note that the XPS has a great many features we are not interested in. He created CSS gui's for the harp scan, collimator, expert motor control, etc.
With luck Nerses will arrive by early Sep, all depends on satisfying new DOE rules for visitors.
Good progress wiring the Rack Room, should be done by the end of the month. Then the racks will be moved in and bolted into place. At this point the Networking group will install fibers between the Telecom Room and the Rack Room, including extras for Accelerator. Immediately after they can install our main switch and enable our networks. At this point we'll remove temporary networking to the Telecom Room and start using our networks, e.g. for the Solenoid test.
- FT schedule being updated by Tom, then Elliott will tie everything in to Tom's infrastructure activities.
- Looks like original schedule delayed by 2 weeks to a month with Tom's update.
- Still not clear if the cryo can will be delivered by the end of Aug.
- George almost done with FMEA.
- Josh will produce OSP, THA, etc. for the test.
- Giles produced a draft OSP for power supply maintenance.
- Power supply now in final location, will be hooked up soon.
- Work on power supply should be complete by the end of Aug.
- Vacuum system designed, parts on order.
- We ordered more turbo pumps.
- Quench detector mods waiting for George to determine max voltage during quench. Chris knows where to get the parts.
- Dave B very busy with FDC but expects to have time to complete Danfysik communications rewrite.
- Josh noted that A-B micro-logix systems may be interesting due to low overhead and module costs. Hovanes pointed out that there should be a good reason to use it since it is a new type of hardware with a new way of programming.
- Elliott is finishing up implementing full power, etc. hierarchy support.
- The base system is complete and will not change, so feel free to add to the database.
- Hierarchy support is in separate tables and in the gui, no effect on base tables.
- Hovanes is still evaluating barcode scanners, with luck will have a PR out in a few weeks.
- Josh wants to start labeling things.
- Long discussion about conventions for tags, short names, etc. Details TBD.
- Theo should have a prototype ready for the Solenoid test.
- Most (all?) hall requirements should be met.
- Most important is the ability to edit and comment on entries.
- History of modifications to entries will also be kept, for those who don't like allowing modifications.
- Solenoid test will result in a full evaluation.